# SeaMo: A Multi-Seasonal and Multimodal Remote Sensing Foundation Model

### v1

@@ 4.5.2 Ablation studies of Temporal-Multimodal blocks
We also conducted ablation studies on the design of Temporal-Multimodal (TM) blocks. Table 11 presents our results. The model without TM blocks performs worse than the model with TM blocks, as also evidenced in Table 12. We explored two designs within the TM blocks: “decouple” and “fuse”.
As depicted in Figure 14, in the “fuse” design, a linear layer is used to fuse tokens from previous seasons with tokens from the current season. In this case, only one set of cross-attention layers is used. On the other hand, in the “decouple” design, tokens from previous seasons are not fused with tokens from the current season. Instead, two sets of independent cross-attention layers are used to interact with the modality tokens at the current season. Analyzing the results in Table 11, it is noteworthy that both the “decouple” and “fuse” designs perform better on some downstream tasks. However, considering that the “decouple” design requires more cross-attention layers, resulting in more model parameters and higher data dimensionality during TM blocks, as well as increased computational resources and longer training times, we choose the “fuse” design as the default setting.
Table 12: Ablation studies of pretraining strategy.
@@
Table 13: Ablation Analysis of the proposed SeaMo in various aspects. Highlighted entries represent the best results.
@@ 4.5.4 Model design ablations
(5) Pretraining length: Figure 15 illustrates how pretraining length affects performance across three downstream tasks. It is evident that longer pretraining durations improve model performance, though the rate of improvement gradually diminishes. These ablation studies provide critical insights into the specific aspects of SeaMo that most significantly influence its effectiveness.
@@ 4.5.5 Generalizability test of the pretrained model
One distinguishing feature of RS images compared to natural images is the diversity in data types. For instance, spectral data captured by the same satellite can vary in image dimensions and the number of spectral bands depending on the preprocessing applied. Given that our model was pretrained on spectral images with dimensions of $128\times 128\times 12$, we aimed to assess its sensitivity to variations in image size and the number of spectral bands in downstream tasks. As described in Section 3, data entering the pretraining model must first pass through a modality-specific tokenizer to compress information. The weights of this tokenizer may not be suitable for different types of data. In our design, we opt for positional encoding interpolation for data of varying sizes, while for data with different channel counts, we introduce a new tokenizer. Although this approach does not utilize the original tokenizer’s pretrained weights, the features are still extracted through the pretrained backbone after projecting. Table 14 displays the performance of our pretrained model after fine-tuning on data with different image sizes and spectral band counts. It is evident that while image size has minimal impact on model performance, a reduction in the number of bands significantly decreases performance. Therefore, future work will need to focus on designing a universal tokenizer that accommodates the diverse range of remote sensing data [42].
Table 14: Generalizability test of the model across different data types in the EuroSAT dataset.
@@
4.5.6 Compute
We estimate the computational resources required for both pretraining and fine-tuning our model, SeaMo, with the metrics presented in terms of RTX4090-24GB GPU hours. Detailed estimates are provided in Tab. 15.
Table 15: Estimated GPU hours used for pretraining and finetuning SeaMo.
@@ 5 Conclusion
In this paper, we introduce SeaMo, a foundation model for remote sensing (RS), pretrained in a self-supervised manner using a rich array of multi-seasonal and multimodal data. SeaMo is equipped with feature-extracting encoders, temporal-multimodal fusion blocks (TM blocks), and image-reconstruction decoders. Under a progressive masked image modeling framework, SeaMo utilizes space-unaligned heterogeneous satellite images to learn rich multimodal representations and leverage unimodal multi-seasonal satellite images to learn robust time-invarient spatial representations. SeaMo has demonstrated outstanding performance across a variety of downstream tasks, successfully handling several tasks across optical and SAR datasets, showing great potential for broader RS applications. Key insights from our work include: (1) The design of the foundation model architecture is crucial for learning the relationships between multiple RS properties, such as temporal-multimodal and spatial-temporal properties. (2) A progressive self-supervised pretraining approach, which gradually increases data complexity and refines training strategies, enables the model to develop robust representations at various stages. Looking ahead, we plan to expand our dataset by collecting more multimodal and multi-source data and to scale up the model to ViT-Large and ViT-Huge configurations. Additionally, we aim to enhance SeaMo’s performance by exploring more advanced tokenizer strategies and positional embeddings. We hope this work offers valuable insights into the development of foundation models for RS imagery.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 42271350 and also supported by the International Partnership Program of the Chinese Academy of Sciences under Grant No.313GJHZ2023066FN. The numerical calculations in this study were carried out on the ORISE Supercomputer.

### v2

@@ 4.5.2 Ablation studies of Temporal-Multimodal blocks
We also conducted ablation studies on the design of TM blocks, with the results presented in Table 11. As shown, the model without TM blocks performs worse than the model that incorporates them, as further evidenced in Table 12. Within the TM blocks, we compared two designs: Decouple and Fuse. As illustrated in Figure 14, the Fuse design employs a linear layer to merge tokens from previous seasons with tokens from the current season, using a single set of cross-attention layers. In contrast, the Decouple design keeps tokens from previous seasons separate, interacting with the current season’s modality tokens via two independent sets of cross-attention layers.
Table 11 indicates that both the Decouple and Fuse designs yield improvements on certain downstream tasks. However, given that the Decouple design incurs a higher computational cost due to an increased number of cross-attention layers, more model parameters, and higher data dimensionality during TM block processing, we adopt the Fuse design as the default setting.
Table 12: Ablation studies of pretraining strategy. In this table, colored cells indicate performance changes relative to the baseline: green cells    and    denote improvements, whereas red/pink cells   ,   ,    and    indicate degradations.
@@
Table 13: Ablation analysis of the proposed SeaMo for various aspects. Best results are shown in bold and highlighted in green   .
@@
4.5.3 Ablations studies of temporal length
Table 14 demonstrates the impact of temporal length during the pre-training phase on model performance. The temporal length of 1 means that only multimodal data is used for training, without including temporal data or the TM Block. It is obvious that as the time series increases, the model can learn more robust temporal and spatial representations, which is reflected in the improved performance on downstream tasks. This proves the necessity of our dedicated research and design for temporal information.
Table 14: Ablation studies on temporal length. Best results are shown in bold and highlighted in green   .
BEN EuroSAT BEN-SAR OCSD
Temporal Length mAP Top1 Acc. Top1 Acc. F1 Score
1 87.81 99.30 81.24 52.10
2 87.92 99.32 81.59 53.17
3 88.31 99.33 82.16 53.98
4 (Default) 88.54 99.37 82.23 54.54
@@ 4.5.5 Model design ablations
(5) Pretraining length: Figure 15 illustrates how pretraining length affects performance on three downstream tasks. Longer pretraining durations improve model performance, though the rate of improvement gradually diminishes. These ablation studies provide critical insights into the specific aspects of SeaMo that most significantly influence its effectiveness.
@@ 4.5.6 Generalizability test of the pretrained model
One distinguishing feature of RS images compared to natural images is the diversity in data types. For instance, spectral data captured by the same satellite can vary in image dimensions and the number of spectral bands depending on the preprocessing applied. Given that our model was pretrained on spectral images with dimensions of $128\times 128\times 12$, we aimed to assess its sensitivity to variations in image size and the number of spectral bands in downstream tasks. As described in Section 3, data entering the pretraining model must first pass through a modality-specific tokenizer to compress information. The weights of this tokenizer may not be suitable for different types of data. In our design, we opt for positional encoding interpolation for data of varying sizes, while for data with different channel counts, we introduce a new tokenizer. Although this approach does not utilize the original tokenizer’s pretrained weights, the features are still extracted through the pretrained backbone after projecting. Table 15 displays the performance of our pretrained model after fine-tuning data with different image sizes and spectral band counts. It is evident that while image size has minimal impact on model performance, a reduction in the number of bands significantly decreases performance. Therefore, future work will need to focus on designing a universal tokenizer that accommodates the diverse range of remote sensing data [42].
Table 15: Generalizability test of the model across different data types in the EuroSAT dataset. Best results are shown in bold and highlighted in green   .
@@
4.5.7 Computational burden
We estimate the computational resources required for both pretraining and fine-tuning our model, SeaMo, with the metrics presented in terms of RTX4090-24GB GPU hours. Detailed estimates are reported in Table 16.
Table 16: Estimated GPU hours used for pretraining and fine-tuning SeaMo.
@@ 5 Conclusion
In this paper, we introduce SeaMo, a remote sensing foundation model pretrained in a self-supervised manner using a rich array of multi-seasonal and multimodal data. SeaMo is equipped with feature-extracting encoders, TM blocks, and image-reconstruction decoders. Under a progressive MIM framework, SeaMo utilizes space-unaligned heterogeneous satellite images to learn rich multimodal representations and leverages unimodal multi-seasonal satellite images to learn robust time-invariant spatial representations. SeaMo has demonstrated outstanding performance on a variety of downstream tasks, successfully handling several tasks on optical and SAR datasets, showing great potential for broader RS applications. Key insights from our work include: (1) The design of the foundation model architecture is crucial for learning the relationships between multiple RS properties, such as temporal-multimodal and spatial-temporal properties. (2) A progressive self-supervised pretraining approach, which gradually increases data complexity and refines training strategies, enables the model to develop robust representations at various stages.
Looking ahead, we plan to expand our dataset by collecting more multimodal and multi-source data and to scale up the model to ViT-Large and ViT-Huge configurations. Additionally, we aim to enhance SeaMo’s performance by exploring more advanced tokenizer strategies and positional embeddings. We hope this work will offer valuable insights into the development of foundation models for RS imagery.
Data availability
The SSL4EO-S12 data that support the findings of this study are openly available in https://mediatum.ub.tum.de/1660427 at https://doi.org/10.1109/MGRS.2023.3281651.
The EuroSAT data that support the findings of this study are openly available in https://zenodo.org/records/7711810#.ZAm3k-zMKEA at https://doi.org/10.1109/IGARSS.2018.8519248.
The BigEarthNet data that support the findings of this study are openly available in https://bigearth.net/ at https://doi.org/10.1109/MGRS.2021.3089174.
The fMoW-Sentinel data that support the findings of this study are openly available in https://purl.stanford.edu/vg497cb6002 at https://dl.acm.org/doi/10.5555/3600270.3600285.
The OSCD data that support the findings of this study are openly available in https://ieee-dataport.org/open-access/oscd-onera-satellite-change-detection at https://doi.org/10.1109/IGARSS.2018.8518015.
The SegMunich data that support the findings of this study are openly available in https://huggingface.co/datasets/Moonboy12138/SegMunich/blob/main/TUM_128.zip at https://doi.org/10.1109/TPAMI.2024.3362475.
The DFC2020 data that support the findings of this study are openly available in https://ieee-dataport.org/competitions/2020-ieee-grss-data-fusion-contest at https://doi.org/10.1109/JSTARS.2021.3063849.
The EuroSAT-SAR data that support the findings of this study are openly available in https://huggingface.co/datasets/wangyi111/EuroSAT-SAR at https://doi.org/10.48550/arXiv.2310.18653.
The GLH-Water data that support the findings of this study are openly available in https://jack-bo1220.github.io/project/GLH-water.html#gLH-water-Dataset at https://doi.org/10.1609/aaai.v38i20.30226.
The S1S2-Water data that support the findings of this study are openly available in https://zenodo.org/records/11278238 at https://dx.doi.org/10.1109/JSTARS.2023.3333969.
The real data that support the findings of this study are available from the first author upon reasonable request.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 42271350 and also supported by the International Partnership Program of the Chinese Academy of Sciences under Grant No.313GJHZ2023066FN.